OpenStudy (anonymous):

solve for x and y? 2y-7x=-9 4x+6=y

OpenStudy (anonymous):

imma help you ok?? @27kriss

OpenStudy (anonymous):

implifying 2y + -3x + 5 = 4x + -6y + -2 Reorder the terms: 5 + -3x + 2y = 4x + -6y + -2 Reorder the terms: 5 + -3x + 2y = -2 + 4x + -6y Solving 5 + -3x + 2y = -2 + 4x + -6y Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. 5 + -3x + -4x + 2y = -2 + 4x + -4x + -6y Combine like terms: -3x + -4x = -7x 5 + -7x + 2y = -2 + 4x + -4x + -6y Combine like terms: 4x + -4x = 0 5 + -7x + 2y = -2 + 0 + -6y 5 + -7x + 2y = -2 + -6y Add '-5' to each side of the equation. 5 + -7x + -5 + 2y = -2 + -5 + -6y Reorder the terms: 5 + -5 + -7x + 2y = -2 + -5 + -6y Combine like terms: 5 + -5 = 0 0 + -7x + 2y = -2 + -5 + -6y -7x + 2y = -2 + -5 + -6y Combine like terms: -2 + -5 = -7 -7x + 2y = -7 + -6y Add '-2y' to each side of the equation. -7x + 2y + -2y = -7 + -6y + -2y Combine like terms: 2y + -2y = 0 -7x + 0 = -7 + -6y + -2y -7x = -7 + -6y + -2y Combine like terms: -6y + -2y = -8y -7x = -7 + -8y Divide each side by '-7'. x = 1 + 1.142857143y Simplifying x = 1 + 1.142857143y here your answer hope it helps @27kriss

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

welcome @27kriss